OpenStudy (rayman5766):

What is the area of the triangle? A. 384 cm² B. 240 cm² C. 192 cm² D. 96 cm²

OpenStudy (igreen):

Picture?

OpenStudy (igreen):

\(A = \dfrac{1}{2} b h\) \(A = \dfrac{1}{2}(24)(16)

OpenStudy (rayman5766):

its hard trying to type all of this in ok so give me a minuet

OpenStudy (igreen):

\(A = \dfrac{1}{2}(24)(16)\)

OpenStudy (igreen):

Multiply those three

OpenStudy (rayman5766):

ell not hard but you know

OpenStudy (rayman5766):

192 c

OpenStudy (igreen):

Yep, that's our answer.